OpenStudy (anonymous):

Factor the polynomial completely. z^3(y^2 − 9) + 125(y^2 − 9)

OpenStudy (anonymous):

(z^3+125)(y^2-9)

OpenStudy (anonymous):

for y^2-9 = (y+3)(y-3)

OpenStudy (anonymous):

For z^3+125 = z^3 +5^3 = (z+5)(z^2-5z+25)

OpenStudy (anonymous):

(z+5)(z^2-5z+25)(y+3)(y-3)